# ReCLIP++: Learn to Rectify the Bias of CLIP for Unsupervised Semantic Segmentation

### v2

@@ 1 Introduction
Semantic segmentation aims to attach a semantic label to each pixel of an image. Since the rise of deep learning [1, 2, 3, 4, 5], semantic segmentation has been widely adopted in real-world applications, e.g., autonomous driving, medical image segmentation, etc. Conventional approaches [6, 7, 8, 9] for semantic segmentation have achieved remarkable performance. However, the superior performance of those methods relies heavily on large amounts of fully annotated masks. Collecting such high-quality pixel-level annotations can be both time-consuming and expensive, e.g., some annotations for specialized tasks require massive expert knowledge, some are even inaccessible due to privacy reasons, etc. Therefore, it is necessary to explore unsupervised semantic segmentation where only images without annotations are available.
@@
We conduct extensive experiments on five standard semantic segmentation benchmarks, including PASCAL VOC [25], PASCAL Context [26], ADE20K [27], Cityscapes [28] and COCO Stuff [29]. Experiment results demonstrate that ReCLIP++ performs favorably against previous state-of-the-arts. Notably, on PASCAL VOC, our method outperforms MaskCLIP+ [20] by $15.4$% and CLIP-S4 [23] by $13.4$% mIoU. Extensive ablation studies verify the effectiveness of each design in our framework.
@@ 2 Related Work
Pre-trained vision-language models. Pre-trained vision-language models (VLMs) [30, 31, 32, 33, 34] have developed rapidly with the help of large-scale image-text pairs available on the Internet. Recently, CLIP [19], ALIGN [35] and Slip [36] have made great progress on learning visual and textual representations jointly by using contrastive learning. With the image-level alignment with text, pre-trained VLMs have a strong ability for zero-shot classification task and can be transferred to various downstream tasks, such as object detection [37, 38] and semantic segmentation [20, 39, 40].
@@ 3.1 Baseline: Directly Segment with CLIP
Each text $Q_{i}$ in Query $Q=\{Q_{1},Q_{2},\cdots,Q_{C}\}$ ($C$ is the number of classes) is an ensembling of several manually designed templates, e.g., “a good/large/bad photo of a $[CLS]$”, where $[CLS]$ denotes a specific class name. Passing $Q_{i}$ through the text encoder, we obtain different text embeddings for different manually designed templates. We average all those text embeddings as the embedding of $Q_{i}$. Then the embeddings of all the $Q_{i}$ in $Q$ make up the query text feature $W_{q}\in\mathbb{R}^{C\times D}$. We treat text feature $W_{q}$ as the weight of the segmentation head to perform 1 $\times$ 1 convolution. By sending feature $Z$ to the segmentation head, we get a Query logit map $M_{q}\in\mathbb{R}^{n\times C}$. Then the segmentation mask can be predicted by $\arg\max$ operation on $M_{q}$ across the class channel $C$.
@@ 3.2 Bias Extraction
where each $[v_{l}](l\in\{1,...,L\})$ is a vector with dimension $D$ and serves as a word embedding. The $L$ is a constant representing the number of word embeddings to learn. Totally, there are $77$ word embeddings for a textual input of CLIP. Among these, two word embeddings are used for representing class names and two for indicating the start and end of a textual input. Thus, we set $L$ to $73$.
@@ 3.3 Rectified Mask Generation
Though the rectified logit map $M$ can be directly used to generate semantic masks, the local context information is not considered, which may result in non-smooth mask predictions. Thus, different from our conference version [24], we additionally introduce a mask decoder $\mathcal{F}_{dec}$ to transform $M$ into a smoother and more contextual logit map. In our implementation, $\mathcal{F}_{dec}$ consists of a convolutional layer with a $5\times 5$ kernel and a batch normalization layer. We concatenate the rectified logit map $M$ to the visual feature $Z$ of CLIP and pass them through the mask decoder $\mathcal{F}_{dec}$ to generate the rectified output $M_{o}\in\mathbb{R}^{n\times C}$, i.e.,
@@ 3.4 Bias Rectification via Contrastive Loss
In ReCLIP [24], we directly calculate the similarity scores between the class-level text features of manually designed prompts, e.g., “a photo of a $[CLS]$”, and the image-level visual feature extracted by CLIP. We select the set of classes whose scores are higher than a threshold as the hypothesis. As CLIP is pre-trained to align image-level visual features with text features, it may focus on the most salient objects in an image and fail to recognize every existing object, rendering the multi-label hypothesis generated for ReCLIP less accurate.
@@
=$\mathcal{H}=\{k|k\in\{1,2,...,C\}\text{ and }f(k)>t\}.$ (6)
Bias Rectification via Contrastive Loss. We then design a contrastive loss to supervise the rectification process via aligning masked features with text features of corresponding classes. Specifically, we apply $\hat{M}$ to the feature $Z$ and perform a global average pooling to get the class-level masked features $Z_{g}\in\mathbb{R}^{C\times D}$, which encode the features of regions belonging to different classes. We compute the similarities between $Z_{g}$ and text features of all the categories. As the masked features $Z_{g}$ represent the objects of interest without context, we utilize $W_{q}$ generated by text input $Q$ for similarity computation. With generated $\mathcal{H}$, the contrastive loss is computed as
@@
where $S_{k,j}$ denotes the cosine similarity between visual feature of class $k\in\mathcal{H}$ and the text feature of $j$-th ($j\in\{1,2,\cdots,C\}$ category and the $\tau$ is a constant.
@@ 4.1 Setup
4.2 Comparison with previous SOTA
[…]
@@ 4.2 Comparison with previous SOTA
GroupViT [57] CVPR’22 TL-OVSS $79.7$ $23.4$ $9.2$ $11.1$ $15.3$
@@
CLIP [19] ICML’21 TF-OVSS $41.8$ $9.2$ $2.1$ $5.5$ $4.4$
@@
ClearCLIP [63] ECCV’24 $81.9$ $34.9$ $16.2$ $18.3$ $22.9$
MaskCLIP+ ECCV’22 C-USS $70.0^{\dagger}$ $31.1$ $12.2^{\dagger}$ $25.2^{\dagger}$ $19.5^{\dagger}$
CLIPpy [22] ICCV’23 $54.6$ $13.5$ $22.3$
ReCo [21] NeurIPS’22 $55.2^{\dagger}$ $26.2$ $19.3$ $26.3$
ReCo [21](w/o. CE) NeurIPS’22 $54.8$ $23.1$ $18.2$ $20.8$
CLIP-S4 [23] CVPR’23 $72.0$ $33.6$
@@
Comparison. The comparisons with previous state-of-the-art methods on five benchmarks are demonstrated in Table 1. From Table 1, we have the following observations: (1) C-USS methods consider the same setting as our method. We observe that ReCLIP++ outperforms previous state-of-the-art C-USS methods obviously, achieving new state-of-the-arts on all five benchmarks. Since ReCo [21] employs a “context elimination” (CE) trick which introduces prior knowledge to assist the training, we also report the results of ReCo by removing this trick (ReCo w/o. CE in Table 1) for a fair comparison. Notably, ReCLIP++ outperforms MaskCLIP+ by $15.4$%, $5.0$%, $4.2$%, $1.3$% and $4.3$% respectively on the five datasets and outperforms CLIP-S4 by $13.4$% and $2.5$% on PASCAL VOC and PASCAL Context respectively. All these results verify the effectiveness of our ReCLIP++ in rectifying the bias of CLIP to assist unsupervised semantic segmentation (2) ReCLIP++ outperforms previous state-of-the-art TF-OVSS methods, implying that the bias of CLIP in complex visual understanding tasks cannot be fully rectified by simply modifying its architecture without any training. (3) TL-OVSS methods consider a different setting and are trained with large-scale image-caption pairs. Strictly speaking, our method cannot be directly compared with those works. However, the superior performance of ReCLIP++ compared to those typical TL-OVSS methods still demonstrates the effectiveness of ReCLIP++.
@@ 4.3 Comparison with Conference Version
In Table 2, we verify the effectiveness of each technical improvement of ReCLIP++ beyond the rectification stage of ReCLIP (rec.) [24] (our conference version): 1) we additionally introduce a decoder that takes the rectified logit map and visual feature of CLIP as input and outputs smoother and more contextual masks (denoted as “Decoder”); 2) we design a new strategy to generate a more accurate label hypothesis for each image (denoted as “Label Hypothesis”); 3) we optimize the design of the bias encoding scheme to independently encode the class-preference and space-preference bias (denoted as “Independent”). In Table 2, we observe that when sequentially adding each technical component, the segmentation performance is consistently improved, which verifies the effectiveness of each technical contribution.
@@
Table 2: Technical improvements of ReCLIP++ compared with ReCLIP. The first line is the results of ReCLIP after the rectification stage and the following lines are ablations experiments on the technical improvements of ReCLIP++ beyond ReCLIP. Results show that each newly added technical improvement of ReCLIP++ contributes to better results than ReCLIP.
ReCLIP (rec.) +Decoder +Label Hypothesis +Independent PASCAL VOC PASCAL Context
@@
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\bf{85.4}$ $\bf{36.1}$
@@ 4.4 Ablation study
Query Reference VOC Context Cityscapes
@@
Subtraction Addition VOC Context Cityscapes
@@
Figure 3: Sensitivity to $t$ and $r$ in the image-level multi-label hypothesis generation. (a) Sensitivity to threshold $t$. (b) Sensitivity to $r$.
@@
Setting VOC Context Cityscapes
@@
Class-preference Space-preference VOC Context
@@
Bias Rectification Evaluation. In this ablation, we further evaluate the rectification effect of two biases both qualitatively and quantitatively.
@@
Method Class-preference Space-preference
@@
Table 8: Distillation results of ReCLIP++. Results show that distillation harms the segmentation performance.
Method VOC Context ADE20K

### v3

@@ 1 Introduction
Semantic segmentation aims to attach a semantic label to each pixel of an image. Since the rise of deep learning [1, 2, 3, 4, 5, 6], semantic segmentation has been widely adopted in real-world applications, e.g., autonomous driving, medical image segmentation, etc. Conventional approaches [7, 8, 9, 10] for semantic segmentation have achieved remarkable performance. However, the superior performance of those methods relies heavily on large amounts of fully annotated masks. Collecting such high-quality pixel-level annotations can be both time-consuming and expensive, e.g., some annotations for specialized tasks require massive expert knowledge, some are even inaccessible due to privacy reasons, etc. Therefore, it is necessary to explore unsupervised semantic segmentation where only images without annotations are available.
@@
We conduct extensive experiments on five standard semantic segmentation benchmarks, including PASCAL VOC [26], PASCAL Context [27], ADE20K [28], Cityscapes [29] and COCO Stuff [30]. Experiment results demonstrate that ReCLIP++ performs favorably against previous state-of-the-arts. Notably, on PASCAL VOC, our method outperforms MaskCLIP+ [21] by 15.415.4% and CLIP-S4 [24] by 13.413.4% mIoU. Extensive ablation studies verify the effectiveness of each design in our framework.
@@ 2 Related Work
Pre-trained vision-language models. Pre-trained vision-language models (VLMs) [31, 32, 33, 34, 35] have developed rapidly, largely due to the availability of large-scale image-text pairs available on the Internet. Early VLMs typically rely on joint embedding spaces and multi-modal transformers to align visual and textual information. They often use objectives such as masked language modeling, image-text matching, and cross-modal retrieval. More recently, contrastive learning-based models such as CLIP [20], ALIGN [36], and SLIP [37] have achieved remarkable progress by training on hundreds of millions of image-text pairs. These models employ a dual-encoder architecture, where images and texts are encoded separately, and a contrastive loss is used to maximize the similarity of matched image-text pairs while minimizing the similarity of mismatched pairs. This approach enables efficient, large-scale training and robust alignment at the image level. As a result, pre-trained VLMs demonstrate strong zero-shot transfer capabilities, excelling in tasks such as zero-shot classification, and can be adapted to various downstream applications, including object detection [38, 39] and semantic segmentation [21, 40, 41].
@@ 3.1 Baseline: Directly Segment with CLIP
Each text $Q_{i}$ in Query $Q=\{Q_{1},Q_{2},\cdots,Q_{C}\}$ ($C$ is the number of classes) is an ensembling of several manually designed templates, e.g., “a good/large/bad photo of a [CLS][CLS]”, where $[CLS]$ denotes a specific class name. Passing $Q_{i}$ through the text encoder, we obtain different text embeddings for different manually designed templates. We average all those text embeddings as the embedding of $Q_{i}$. Then the embeddings of all the $Q_{i}$ in $Q$ make up the query text feature $W_{q}\in\mathbb{R}^{C\times D}$. We treat text feature $W_{q}$ as the weight of the segmentation head to perform 1 $\times$ 1 convolution. By sending feature $Z$ to the segmentation head, we get a Query logit map $M_{q}\in\mathbb{R}^{n\times C}$. Then the segmentation mask can be predicted by $\arg\max$ operation on $M_{q}$ across the class channel $C$.
@@ 3.2 Bias Extraction
where each $v_{l}](l\in\{1,...,L\})$ is a vector with dimension $D$ and serves as a word embedding. The $L$ is a constant representing the number of word embeddings to learn. Totally, there are $77$ word embeddings for a textual input of CLIP. Among these, two word embeddings are used for representing class names and two for indicating the start and end of a textual input. Thus, we set $L$ to $73$.
@@ 3.3 Rectified Mask Generation
Though the rectified logit map $M$ can be directly used to generate semantic masks, the local context is not fully considered, which may result in non-smooth mask predictions. The reason is as follows. The query logit map $M_{q}$ is computed as the cosine similarity between each patch feature and textual features, treating each patch independently without considering information from neighboring patches. The bias logit map $M_{b}$ is generated from the reference feature $W_{r}$ (which encodes only class-related information and is spatially invariant) and the positional feature $W_{p}$ (which is encoded with $1\times 1$ convolution and does not contain semantic information), without considering spatial interaction. Since both $M_{q}$ and $M_{b}$ are computed independently for each patch, the resulting rectified logit map $M$ is constructed in a patch-wise manner. As a result, directly using the rectified logit map $M$ for mask prediction may result in non-smooth masks because each patch is predicted without awareness of its local neighbors.
Thus, different from our conference version [25], we additionally introduce a mask decoder $\mathcal{F}_{dec}$ to transform $M$ into a smoother and more contextual logit map. In our implementation, $\mathcal{F}_{dec}$ consists of a convolutional layer with a $5\times 5$ kernel and a batch normalization layer. We concatenate the rectified logit map $M$ to the visual feature $Z$ of CLIP and pass them through the mask decoder $\mathcal{F}_{dec}$ to generate the rectified output $M_{o}\in\mathbb{R}^{n\times C}$, i.e.,
@@ 3.4 Bias Rectification via Contrastive Loss
In ReCLIP [25], we directly calculate the similarity scores between the class-level text features of manually designed prompts, e.g., “a photo of a [CLS][CLS]”, and the image-level visual feature extracted by CLIP. We select the set of classes whose scores are higher than a threshold as the hypothesis. However, the objects in an image may have different scales (large/small) and may be located arbitrarily (center/boundary). As CLIP is pre-trained to align image-level visual features with text features, it may focus on the most salient objects in an image and fail to recognize every existing object. Thus, generating the multi-label hypothesis based on the image-level visual features (like ReCLIP) cannot fully discover the underlying categories in an image, especially for categories whose objects are usually small or less salient in an image.
@@
$\mathcal{H}=\{k|k\in\{1,2,...,C\}\text{ and }f(k)>t\}.$ (6)
The rationale behind the strategy of multi-label hypothesis in ReCLIP++ is that the objects that are not salient globally may be salient locally. By aggregating the predictions from multiple local regions, ReCLIP++ can better discover existing categories in an image, especially for those that are small or less salient.
Bias Rectification via Contrastive Loss. We then design a contrastive loss to supervise the rectification process via aligning masked features with text features of corresponding classes. Specifically, we expand $\hat{M}$ to $\hat{M^{\prime}}\in\mathbb{R}^{n\times C\times D}$ and $Z$ to $Z^{\prime}\in\mathbb{R}^{n\times C\times D}$, apply $\hat{M^{\prime}}$ to the feature $Z^{\prime}$, and perform a global average pooling to get the class-level masked features $Z_{g}\in\mathbb{R}^{C\times D}$, which encode the features of regions belonging to different classes:
$Z_{g}=\text{Avg-Pool}(\hat{M^{\prime}}\otimes Z^{\prime}),$ (7)
where $\otimes$ represents element-wise multiplication.
We compute the similarities between $Z_{g}$ and the text features of all the categories. As the masked features $Z_{g}$ represent the objects of interest without context, we utilize $W_{q}$ generated by text input $Q$ for similarity computation. With generated $\mathcal{H}$, the contrastive loss is computed as
@@
where $S_{k,j}$ denotes the cosine similarity between visual feature of class $k\in\mathcal{H}$ (denoted as $Z_{g,k}$) and the text feature of $j$-th ($j\in\{1,2,\cdots,C\}$ category (denoted as $W_{q,j}$) and the $\tau$ is a constant. Among it, $S_{k,j}$ can be calculated as follows ($\cdot$$\cdot$> means dot product):
$S_{k,j}=\frac{<Z_{g,k},W_{q,j}>}{\|Z_{g,k}\|\|W_{q,j}\|}.$ (9)
@@ 4.1 Setup
4.2 Comparison with previous state-of-the-arts
Baselines. We mainly compare our method with three types of semantic segmentation methods to verify the superiority of our method: (1) Trainable language-guided methods for OVSS task (TL-OVSS), including GroupViT [58], CoCu [63], and TCL [40]; (2) CLIP-based methods for training-free OVSS task (TF-OVSS), including CLIP [20] (vanilla CLIP visual encoder), MaskCLIP [21] (CLIP visual encoder with modified last-attention block), SCLIP [65], and ClearCLIP [64]; (3) CLIP-based methods for USS task (C-USS), including MaskCLIP+ [21], CLIPpy [23], ReCo [22], CLIP-S4 [24] and ReCLIP [25]. We directly cite the corresponding results from the original papers, except that $\dagger$} means the results are obtained by running the officially released source code and $\ddagger$} means the results are cited from TCL [40]. All the numbers reported are presented as percentages. Among these, TL-OVSS methods rely on weak annotations like image-caption pairs to train the model, while TF-OVSS methods can directly perform open-vocabulary segmentation without any training. C-USS methods share the same language-guided unsupervised setting with ReCLIP++.
Table 1: Comparison with trainable language-guided methods for OVSS (TL-OVSS), CLIP-based methods for training-free OVSS (TF-OVSS), and CLIP-based methods for USS (C-USS) on five various benchmarks. † means the results are obtained by running the officially released source code and $\ddagger$} means the results are cited from TCL [40].
@@ 4.2 Comparison with previous state-of-the-arts
GroupViT [58] CVPR’22
TL-OVSS
$79.7$ $23.4$ $9.2$ $11.1$ $15.3$
@@
CLIP [20] ICML’21
TF-OVSS
$41.8$ $9.2$ $2.1$ $5.5$ $4.4$
@@
ClearCLIP [64] ECCV’24 $80.9$ $34.9$ $16.2$ $18.3$ $22.9$
MaskCLIP+ ECCV’22
C-USS
$70.0^{\dagger}$ $31.1$ $12.2^{\dagger}$ $25.2^{\dagger}$ $19.5^{\dagger}$
CLIPpy [23] ICCV’23 $54.6$ - $13.5$ $22.3$ -
ReCo [22] NeurIPS’22 $55.2^{\dagger}$ $26.2$ - $19.3$ $26.3$
ReCo [22](w/o. CE) NeurIPS’22 $54.8$ $23.1$ - $18.2$ $20.8$
CLIP-S4 [24] CVPR’23 $72.0$ $33.6$ - - -
@@
Comparison. The comparisons with previous state-of-the-art methods on five benchmarks are demonstrated in Table 1. From Table 1, we have the following observations: (1) C-USS methods consider the same setting as our method. We observe that ReCLIP++ outperforms previous state-of-the-art C-USS methods obviously, achieving new state-of-the-arts on all five benchmarks. Since ReCo [22] employs a “context elimination” (CE) trick which introduces prior knowledge to assist the training, we also report the results of ReCo by removing this trick (ReCo w/o. CE in Table 1) for a fair comparison. Notably, ReCLIP++ outperforms MaskCLIP+ by 15.415.4%, 5.05.0%, 4.24.2%, 1.31.3% and 4.34.3% respectively on the five datasets and outperforms CLIP-S4 by 13.413.4% and 2.52.5% on PASCAL VOC and PASCAL Context respectively. All these results verify the effectiveness of our ReCLIP++ in rectifying the bias of CLIP to assist unsupervised semantic segmentation (2) ReCLIP++ outperforms previous state-of-the-art TF-OVSS methods, implying that the bias of CLIP in complex visual understanding tasks cannot be fully rectified by simply modifying its architecture without any training. (3) TL-OVSS methods consider a different setting and are trained with large-scale image-caption pairs. Strictly speaking, our method cannot be directly compared with those works. However, the superior performance of ReCLIP++ compared to those typical TL-OVSS methods still demonstrates the effectiveness of ReCLIP++.
@@ 4.3 Comparison with Conference Version
In Table 2, we verify the effectiveness of each technical improvement of ReCLIP++ beyond the rectification stage of ReCLIP (rec.) [25] (our conference version): 1) we additionally introduce a decoder that takes the rectified logit map and visual feature of CLIP as input and outputs smoother and more contextual masks (denoted as “Decoder”); 2) we design a new strategy to generate a more accurate label hypothesis for each image (denoted as “LH”); 3) we optimize the design of the bias encoding scheme to independently encode the class-preference and space-preference bias (denoted as “Indep.(Multiplication)”). In Table 2, we observe that when sequentially adding each technical component, the segmentation performance is consistently improved, which verifies the effectiveness of each technical contribution.
@@
Table 2: Technical improvements of ReCLIP++ compared with ReCLIP. The first line is the results of ReCLIP after the rectification stage, and the following lines are ablation experiments on the technical improvements of ReCLIP++ beyond ReCLIP. “LH” represents Label Hypothesis, while “Indep.” represents independent encoding of bias. Results show that each newly added technical improvement of ReCLIP++ contributes to better results than ReCLIP.
ReCLIP (rec.) Decoder LH Indep.(Addition) Indep.(Multiplication) VOC Context
@@
$\checkmark$ $\checkmark$ $\checkmark$ $81.1$ $33.3$
$\checkmark$ $\checkmark$ $\checkmark$ $\bf{85.4}$ $\bf{36.1}$
@@ 4.4 Ablation study
Query Reference PASCAL VOC PASCAL Context Cityscapes
@@
Effect of different fusion strategies in bias logits generation. In Table 2, we conduct an ablation comparing two fusion strategies for $W_{r}$ and $W_{p}$: addition (Indep.(Addition)) and matrix multiplication (Indep.(Multiplication)). Comparing Line 3 & 4, we find that even when using the addition fusion strategy of $W_{r}$ and $W_{p}$, ReCLIP++ with the independent bias encoding strategy still remarkably outperforms the model without such a strategy. This demonstrates that our independent bias encoding strategy in ReCLIP++ contributes significantly to the performance improvement. By comparing Line 4 & 5, we find that fusing $W_{r}$ and $W_{p}$ via matrix multiplication yields better results. Based on the results above, we adopt matrix multiplication for fusing $W_{r}$ and $W_{p}$ in ReCLIP++ to achieve the best results.
@@
Subtraction Addition PASCAL VOC PASCAL Context Cityscapes
@@
Figure 3: Sensitivity to $t$ and $r$ in the image-level multi-label hypothesis generation. (a) Sensitivity to threshold $t$. In the image-level multi-label hypothesis generation process, we set the window width and height to $r$ times the width and height of the input image. (b) Sensitivity to $r$. In the image-level multi-label hypothesis generation process, $t$ is the threshold for choosing classes with high frequency within window crops. Results of (a) and (b) verify that ReCLIP++ is insensitive to the choices of $t$ and $r$ and within a vast range of hyperparameter choices, ReCLIP++ remains superior to the solution which employs the multi-label hypothesis strategy of ReCLIP (denoted as “w/o. Crop”).
@@
   Setting    PASCAL VOC    PASCAL Context    Cityscapes
@@
Class-preference Space-preference PASCAL VOC PASCAL Context
@@
Bias Rectification Evaluation. In this ablation, we further evaluate the rectification effect of two biases, both qualitatively and quantitatively.
@@
Method Class-preference Rectification Space-preference Rectification
@@
Effect of different inputs of mask decoder $\mathcal{F}_{dec}$. In this ablation, we design an ablation on the effect of different inputs of the mask decoder $\mathcal{F}_{dec}$ in Table 8. Line 1 refers to using only the rectified logit map $M$ as the input of the decoder, while Line 2 refers to ReCLIP++, which uses both the feature $Z$ and the rectified logit map $M$ as inputs. Results on PASCAL VOC [26] and PASCAL Context [27] show that ReCLIP++ (Line 2) outperforms the strategy that uses only $M$ as the input of $\mathcal{F}_{dec}$ (Line 1). Besides, we quantitatively evaluate both class-preference and space-preference bias rectification effects using our designed metrics described in Section 4.4 (“Bias Rectification Evaluation (2)”). Results in Table 8 show that including $Z$ as the input of $\mathcal{F}_{dec}$ further improves bias rectification compared to using $M$ alone.
Table 8: Ablations on inputs of the decoder. We compare these two strategies by both performance and bias rectification effect. The performance is represented by mIoUs on PASCAL VOC and PASCAL Context and the bias rectification is represented by our designed metrics. A higher number of “Class-Preference Rectification” or “Space-Preference Rectification” means a stronger bias rectification effect.
$Z$ $M$ VOC Context Class-Preference Rectification Space-Preference Rectification
$\checkmark$ $73.6$ $28.0$ $0.34$ 0.35-0.35
$\checkmark$ $\checkmark$ 85.4 36.1 0.40 0.05
@@
Table 9: Distillation results of ReCLIP++. Results show that distillation cannot bring consistent improvement on all benchmarks anymore.
   Method    PASCAL VOC    PASCAL Context    ADE20K